OpenStudy (anonymous):

(7√2)(2√6) in simplest form and i need to show my work

OpenStudy (anonymous):

multiply those , u will get \[14\sqrt{12}\]

OpenStudy (anonymous):

\[\sqrt{12}=2\sqrt{3}\]

OpenStudy (anonymous):

you got it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why? \[14\sqrt{12}=14\times2\sqrt{3}=28\sqrt{3}\]

OpenStudy (anonymous):

ok so then 8√2+3√2=11x2√2=22√2

OpenStudy (anonymous):

no no..you are not multiplying here..just add them

OpenStudy (anonymous):

so 11√4

OpenStudy (anonymous):

yeah,right

OpenStudy (anonymous):

ok what about√50-√32 is it√18

OpenStudy (anonymous):

and 8/√2would be 4

OpenStudy (anonymous):